OpenStudy (anonymous):

16^x=64 solve the equation for x. if necessary round your answer to two decimal places.

OpenStudy (anonymous):

i got 3/2. but it doesnt work.

OpenStudy (anonymous):

\[16^{x}=64\]

OpenStudy (jdoe0001):

keep in mind that \(\bf 16^x= 64 \\ 16 = 2^4, 64 = 2^6\\ 16^x= 64 \implies 2^{4x} = 2^6\)

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

start by rewriting both sides of the equation with the same base... base 4 \[16^x = (4^2)^x = 4^{2x}\] and \[64 = 4^3\] so you are now solving \[4^{2x} = 4^3\] so just equate the powers and solve for x

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

\[4^{2x} is 1048576\]

OpenStudy (anonymous):

\[4^{3} is 64\]

OpenStudy (campbell_st):

so x = 3/2 which is the correct and so go back to the initial question and substitute \[16^{\frac{3}{2}} = 64\]

OpenStudy (anonymous):

thanks.